I'm attempting to assist my daughter with graphing this function/equation or inequality. It's been years since I've done this, I don't even know where to start. Can you please assist?
Here is the problem which needs to be graphed:
d(x)+8-x

Answer by Fombitz(32388)

Did you mean d(x)=8-x??
I hope so.
Pick a range of x values.
Let's choose (-3,3).
Make a table of values,
x=-3,d(-3)=8-(-3)=11
x=-2,d(-2)=8-(-2)=10
x=-1,d(-1)=8-(-1)=9
x=0,d(0)=8-(0)=8
x=1,d(1)=8-(1)=7
x=2,d(2)=8-(2)=6
x=3,d(3)=8-(3)=5
.
.
.

.
.
.
Then connect the points with a line,